Maynard Steel plans to pay a dividend of $ 2.92 this year. The company has an expected earnings growth rate of 3.6 % per year and an equity cost of capital of 9.2 %.

a. Assuming​ Maynard's dividend payout rate and expected growth rate remain​ constant, and that the firm does not issue or repurchase​ shares, estimate​ Maynard's share price. ​(Round to the nearest​ cent.)

b. Suppose Maynard decides to pay a dividend of $ 0.94 this year and use the remaining $ 1.98 per share to repurchase shares. If​ Maynard's total payout rate remains​ constant, estimate​ Maynard's share price. ​(Round to the nearest​ cent.)

c. If Maynard maintains the dividend and total payout rate in ​(b​), at what rate are​ Maynard's dividends and earnings per share expected to​ grow? (Round by two decimals)
Answer #1

a). P0 = D1 / [r - g] = $2.92 / [0.092 - 0.036] = $2.92 / 0.056 = $52.14

b). P0 = Total Payout next year / [r - g] = $2.92 / [0.092 - 0.036] = $2.92 / 0.056 = $52.14

c). Growth Rate = Cost of Capital - Dividend Yield

= 0.092 - [$0.94 / $52.14] = 0.092 - 0.018 = 0.0740, or 7.40%